Hi, i would like some help on this Linear equation:
-3x-3y=3
y=-5x-17
--
My answer ended up being (-4,5), but my friend says the answer is (-4,3). We are both terribly confused ! When we check the equation both answers are correct.

-Thank you

Found 2 solutions by rfer, josgarithmetic:
Answer by rfer(16322)

y=-5x-17
----
-3x-3y=3
-3x-3(-5x-17)=3
-3x+15x+51=3
12x=3-51
12x=-48
x=-4
-------------
-3*-4-3y=3
12-3y=3
-3y=3-12
-3y=-9
y=3
--------------
(-4,3)
---------------
3=5*-4-17
3=20-17
3=3
--------------
5=-5x-17
5+17=-5x
22=-5x
22/-5=x
x=-4.4
Sorry, (-4,5) does not work

Answer by josgarithmetic(39621)

The -3 equation transforms easily to x+y=-1, and the "17" equation becomes -5x-y=17.

Summary of System:
-5x-y=17
x+y=-1

Add both equations: -4x+0=16; therefore, x=-4.
Substitute the value into either equation.
-4+y=-1
y=+4-1
y=+3
Answer: point (-4,3)

When you check, you found both answers as correct?
Let's try the other point you or the friend found, (-4,5):
-5*(-4)-(5)=17?
20-5=17?
15=17? NO.